# Discretion in the Loop: Human Expertise in Algorithm-Assisted College Advising

### v2

@@ A.2 Theoretical Framework
A.2.3 Generalization to Nonbinary Contrasts
Stating three causal conditions of expertly targeted action for nonbinary contrasts requires determining the importance of each causal contrast for each action and algorithmic unobservable valuation. For instance, in determining whether a particular action taken by an advisor (say, recommending a change to a student’s schedule in a meeting), was effective at improving a student’s outcome (say, increasing their chances of graduation), what other action should we contrast it to? Should we contrast it to taking some other action (say, recommending a student speak with their instructor)? If so, which one? Should we contrast it to taking no other action, imagining that the advisor had not met with the student at all? Or should we combine these (and other) possibilities together somehow?
@@ A.2.3 Generalization to Nonbinary Contrasts
A.3 Empirical Analysis of Advisor Expertise
A.3.1 Test Population
@@
A.3.2 Positive vs. Negative Interventions
@@
A.3.3 Test Versions and Parameters
@@
As mentioned in §4.1.2, we greedily pair $L$ “approximately-identical” students at baseline using a Euclidian distance function on fixed-scaled administrative features (excluding their race and ethnicity, which Georgia State University excludes from its predictors), prohibiting cross-semester matches for tests on semesterly outcomes. We select the number of pairs $L=100$ for cumulative tests, $L=800$ for semesterly GPA tests, and $L=700$ for semesterly persistence tests, to yield an acceptable tradeoff of the number of pairs (which increases test power) with the quality of the matches (the more imperfect the matches, the higher the test’s error).
@@
A.3.4 Number of Swaps and Test Power
@@ 20. Develop degree planner 1810 41 70 10 23
Table 7: At $L=800$ pairs for GPA outcomes and $L=700$ for persistence outcomes, the above table records the total number of swaps of intervention $i$ for approximately algorithmically-indistinguishable pairs that could decrease ()($\downarrow$)( ↓ ) vs. increase ()($\uparrow$)( ↑ ) the mean squared error of the intervention at predicting the outcomes. Counts of the number of times the intervention was performed total are also included.
A.3.5 Full Results
@@
Intervention Term GPA (+)(+)( + ) Term GPA (+)(+)( + ) Next-term persistence (+)(+)( + ) Next-term persistence ()(-)( - )
@@
Intervention Persistence (+)(+)( + ) Persistence ()(-)( - ) Graduation (+)(+)( + ) Graduation ()(-)( - )
@@
Intervention Next-term persistence (+)(+)( + ) Next-term persistence ()(-)( - ) Term GPA (+)(+)( + ) Term GPA ()(-)( - )
@@
Intervention Next-term persistence (+)(+)( + ) Next-term persistence ()(-)( - ) Term GPA (+)(+)( + ) Term GPA ()(-)( - )
@@
A.3.6 Correlations Between Interventions
@@
A.4 Exploratory Qualitative Analysis
A.4.1 Codebook for the Exploratory Qualitative Analysis.
In our analysis, codes served as succinct, essence-capturing labels that represent meaningful patterns in the data [61]. We organized codes hierarchically, with high-level categories capturing key differences in advisor style and aims (e.g. discussing a topic open-endedly, making directive suggestions to the student, or building rapport with the student) and finer-grained codes differentiating interactions within these categories (for instance, the discussion topic). We report preliminary frequencies of these applied codes by each advisor in Table 13. Our complete codebook is provided below.
  • Academic Discussion: Do the advisor and student discuss an academic topic (without advisor making a recommendation)?
    • Scheduling - discussing course schedules, registration timelines, or planning for upcoming semesters
    • Academic administrative - addressing administrative issues like degree requirements, deadlines, navigating transfers, or institutional processes
    • Major/minor selection - discussing the student choice of major/minor
    • Academic performance - reviewing grades, progress, or challenges in coursework
    • EAB/algorithmic review - reviewing algorithmically-suggested or institutional benchmarks for the student, like EAB’s Success Markers or Academic Improvement Plan warnings
  • Nonacademic Discussion: Do the advisor and student discuss a non-academic topic (without advisor making a recommendation)?
    • Career - talking about career aspirations, opportunities, career fairs, or preparation
    • Financial aid - discussing applying for or maintaining financial aid packages/status
    • Financial trouble - discussing problems students are having with finances
    • Other nonacademic - discussing any other nonacademic matter, such as extracurriculars, housing/commuting, or family, medical concerns
  • Academic Suggestion: How is the advisor guiding the student academically?
    • Follow up with advisor - the advisor asks students to check in with them, schedule a meeting with them, or follow up in the future with updates about a specifically academic matter
    • Follow up with another person - the advisor asks students to speak with another person about a specifically academic matter
    • Guiding/steering the student - the advisor tells a student how to alter their program of study, for instance by changing major, recommending a schedule, or adding/dropping a class
  • Nonacademic Suggestion: How is the advisor guiding the student non-academically?
    • Follow up with advisor - the advisor asks students to check in with them later on or lets them know that they can check in with future questions about specifically nonacademic matters
    • Other suggestion - the advisor makes any other sort of suggestion about specifically non-academic matters, for instance referring students to a campus office or recommending a student pursue an extracurricular
  • Generic Suggestion: How is the advisor guiding the student across both academic/nonacademic matters?
    • Follow up with advisor - the advisor asks students to check in with them later / follow up later, without mentioning any specific topic
  • Relationship: How does the advisor get to know/support the student emotionally?
    • Empathy/sympathy - expressing concern, understanding, or emotional support for the student’s circumstances or challenges
    • Student affect - observing and noting the student’s emotional state or behavior during the interaction
  • Intervention: Does the advisor take action on behalf of the student?
    • Changing the student’s schedule - directly altering a student’s schedule on their behalf
    • Contacting student’s parents - communicating directly with a student’s family members to address concerns or provide important information
    • Speaking with department/instructor/administrator - The advisor reaches out to a student’s department, instructor, or an administrator on behalf of them
    • Signing / filling out form for student - The advisor signs or completes a form on behalf of the student
Code Category Sub-code Advisor 1 Advisor 2 Advisor 3
Academic Discussion Scheduling 750 337 244
Academic administrative 298 90 28
Major/Minor Selection 214 97 80
Academic performance 190 33 38
EAB/algorithmic review 495 1 4
Academic Suggestion Follow up with advisor 185 13 19
Follow up with another person 52 29 23
Guiding/steering the student 115 51 129
Nonacademic Discussion Career 62 27 50
Financial aid 20 6 8
Financial trouble 33 19 24
Other nonacademic 26 7 32
Nonacademic Suggestion Follow up with advisor 39 0 7
Other suggestion 52 7 20
Generic Suggestion Follow up with advisor 642 1 60
Relationship Empathy/sympathy 42 3 20
Student affect 40 16 94
Intervention Changing student’s schedule 1 0 2
Contacting student’s parents 7 0 1
Speaking with department/instructor 2 2 1
Signing the form for student 24 15 2
@@ A.4 Exploratory Qualitative Analysis
Table 13: Preliminary frequencies of the applied codes by each advisor during student-advising meetings, based on our exploratory qualitative analysis.
A.5 Qualitative Analysis of Advisor Expertise
A.5.1 Codebook for Qualitative Analysis of Non-Algorithmic Information
@@
A.5.2 Comments Mentioning Reasons
@@
Instead of analyzing reasons for taking the intervention for mentions of non-algorithmic information, we analyzed comments for customizations of schedules. See Appendix A.5.3 for details about schedule customization.
A.5.3 Results for Intervention 17 (Recommend Schedule)
@@
A.5.4 Comparison to Insignificant Interventions
Unlike the significant interventions, which incorporate diverse, non-algorithmic information, two of the three insignificant interventions, Intervention 9 (Address Financial Issues) and Intervention 10 (Address Non-financial Collegiate Requirements), are dominated by administrative holds on registration, most commonly due to nonpayment of a balance on the student’s account. These administrative holds are plausibly either predictable from administrative data (for instance, financial holds might be predictable from students’ expected financial contribution) or not outcome-relevant (when the holds are resolved).

### v3

@@ A.2 Theoretical Framework
A.3 Proof of Proposition 4.3
Proof.
By the Markov condition, if $A_{t}\not\perp Y_{t+1}|\ \mathbf{X}_{t}$ then $A_{t}\not\perp_{\mathcal{G}_{\mathcal{M}}}Y_{t+1}|\ \mathbf{X}_{t}$. Thus, by the definition of $d$-separation, either $A_{t}\to Y_{t+1}$ or $\mathbf{U}_{t}\to A_{t}$. First, consider the case that $A_{t}\rightarrow Y_{t+1}$. We will show that in this case, the impactful action criterion is met, which requires showing the following inequality holds:
$\mathbb{P}(Y_{t+1}=1|do(A_{t}=1),X_{t}=x,U_{t}=u)\neq\mathbb{P}(Y_{t+1}=1|do(A_{t}=0),X_{t}=x,U_{t}=u)$ (1)
By Assumption 4.3 of causal minimality, for any $Y$, for all $X\in\mathbf{PA}_{Y}$, $X\not\perp Y|PA_{Y}\backslash\{X\}$ [69, p.109]. Since $A_{t}\in\mathbf{PA}_{Y_{t+1}}$ by assumption, for some $a$, $y$, $x$ and $u$, this implies:
$\mathbb{P}(Y_{t+1}=y,A_{t}=a|X_{t}=x,U_{t}=u)\neq\mathbb{P}(Y_{t+1}=y|X_{t}=x,U_{t}=u)\times\mathbb{P}(A_{t}=a|X_{t}=x,U_{t}=u)$ (2)
Using the product rule, we can manipulate the left hand side of Inequality 2:
$P(Y_{t+1}=y,A_{t}=a|X_{t}=x,U_{t}=u)=P(Y_{t+1}=y|A_{t}=a,X_{t}=x,U_{t}=u)\times P(A_{t}=a|X_{t}=x,U_{t}=u)$
Then cancelling $P(A_{t}=a|X_{t}=x,U_{t}=u)$ from both sides of Inequality 2, we get:
$P(Y_{t+1}=y|A_{t}=a,X_{t}=x,U_{t}=u)$ $\neq\mathbb{P}(Y_{t+1}=y|X_{t}=x,U_{t}=u)$ (3)
Applying the law of total probability, we can now expand the right hand side of Inequality 3 as:
$\mathbb{P}(A_{t}=1|X_{t}=x,U_{t}=u)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)+$
$\mathbb{P}(A_{t}=0|X_{t}=x,U_{t}=u)\times\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$
Without loss of generality, suppose $a=1$. Then by subtraction of $\mathbb{P}(A_{t}=1|X_{t}=x,U_{t}=u)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)$ from both sides of Inequality 3,
$\mathbb{P}(A_{t}=0|X_{t}=x,U_{t}=u)\times\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)\neq$
$\mathbb{P}(A_{t}=0|X_{t}=x,U_{t}=u)\times\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$
And so by cancellation of $\mathbb{P}(A_{t}=0|X_{t}=x,U_{t}=u)$ from both sides,
$\mathbb{P}(Y_{t+1}=y|A_{t}=1,X_{t}=x,U_{t}=u)\neq\mathbb{P}(Y_{t+1}=y|A_{t}=0,X_{t}=x,U_{t}=u)$ (4)
Since $Y_{t+1}$ is binary, we can replace $y$ with 1 (as if $y=0$, this inequality implies the same inequality holds for $Y_{t+1}=1$):
$\mathbb{P}(Y_{t+1}=1|A_{t}=1,X_{t}=x,U_{t}=u)\neq\mathbb{P}(Y_{t+1}=1|A_{t}=0,X_{t}=x,U_{t}=u)$ (5)
Finally, by applying parents do/see [13, p.32] to $A_{t}$ in Inequality 5, we have that Inequality 1 is true, that is, impactful action is met. In the case that $\mathbf{U}_{t}\rightarrow A_{t}$, an analogous proof shows that non-algorithmic action is met. ∎
A.4 Generalization to Nonbinary Contrasts
Stating three causal conditions of expertly targeted action for nonbinary contrasts requires determining the importance of each causal contrast for each action and non-algorithmic information valuation. For instance, in determining whether a particular action taken by an advisor (say, recommending a change to a student’s schedule in a meeting), was effective at improving a student’s outcome (say, increasing their chances of graduation), what other action should we contrast it to? Should we contrast it to taking some other action (say, recommending a student speak with their instructor)? If so, which one? Should we contrast it to taking no other action, imagining that the advisor had not met with the student at all? Or should we combine these (and other) possibilities together somehow?
@@ A.4 Generalization to Nonbinary Contrasts
A.5 Empirical Analysis of Advisor Expertise
A.5.1 Test Population
@@
A.5.2 Positive vs. Negative Interventions
@@
A.5.3 Test Versions and Parameters
@@
As mentioned in §5.1, we greedily pair $L$ “approximately-identical” students at baseline using a Euclidian distance function on fixed-scaled administrative features (excluding their race and ethnicity, which Georgia State University excludes from its predictors), prohibiting cross-semester matches for tests on semesterly outcomes. We select the number of pairs $L=100$ for cumulative tests, $L=800$ for semesterly GPA tests, and $L=700$ for semesterly persistence tests, to yield an acceptable tradeoff of the number of pairs (which increases test power) with the quality of the matches (the more imperfect the matches, the higher the test’s error).
@@
A.5.4 Number of Swaps and Test Power
@@ 20. Develop degree planner 1810 41 70 10 23
Table 8: At $L=800$ pairs for GPA outcomes and $L=700$ for persistence outcomes, the above table records the total number of swaps of intervention $i$ for approximately algorithmically-indistinguishable pairs that could decrease $(\downarrow)$ vs. increase $(\uparrow)$ the mean squared error of the intervention at predicting the outcomes. Counts of the number of times the intervention was performed total are also included.
A.5.5 Full Results
@@
Intervention Term GPA $(+)$ Term GPA $(+)$ Next-term persistence $(+)$ Next-term persistence ()(-)
@@
Intervention Persistence $(+)$ Persistence ()(-) Graduation $(+)$ Graduation ()(-)
@@
Intervention Next-term persistence $(+)$ Next-term persistence ()(-) Term GPA $(+)$ Term GPA ()(-)
@@
Intervention Next-term persistence $(+)$ Next-term persistence ()(-) Term GPA $(+)$ Term GPA ()(-)
@@
1. Seek academic support 162 19 11 11 5
2. Correct schedule 503 55 32 14 11
3. Declare or change major 238 23 18 8 4
4. Progression issues 666 63 35 20 12
5. Adjust course balance 192 22 19 11 10
6. See instructor 73 10 6 3 0
7. Seek academic coaching 14 1 0 0 0
8. Seek departmental or faculty guidance 190 15 23 6 3
9. Address financial issues 338 25 31 9 10
10. Address non-financial collegiate requirements 280 23 26 6 8
11. Seek career counseling and service 307 23 31 5 11
12. Seek engagement opportunities on campus 253 13 29 6 13
13. See professional counselor 28 1 4 2 2
14. Seek disability services on campus 2 0 0 0 0
15. See medical professional 9 1 0 0 0
16. Schedule an appointment 1035 76 45 25 13
17. Recommended schedule 1601 50 66 8 24
18. See primary program advisor 49 6 7 1 3
19. Other (describe in comments) 635 7 19 14 6
20. Academic map/degree planner 1810 41 70 10 23
Table 14: At $L=800$ pairs for GPA outcomes and $L=700$ for persistence outcomes, the above table records the total number of swaps of intervention $i$ for approximately algorithmically-indistinguishable pairs that could decrease $(\downarrow)$ vs. increase $(\uparrow)$ the mean squared error of the intervention at predicting the outcomes. Counts of the number of times the intervention was performed total are also included.
A.5.6 Correlations Between Interventions
@@
Appendix B Exploratory Qualitative Analysis
B.1 Methods
In the first stage of our qualitative analysis, we conduct exploratory coding on all 1,594 advisor meetings comments to understand the nature of advisor-student interactions as documented by those comments. In our analysis, codes served as succinct, essence-capturing labels that represent meaningful patterns in the data [78].
Codes
To determine the codes for our qualitative analysis, Authors 1 and 5 first coded subsamples of comments independently with fine-grained codes, and then met to design a codebook organizing all into three major thematic categories. Author 1 then coded the entire dataset for these three thematic categories using the Atlas.ti software. Our complete codebook with an example quote for each category is provided below.
  • Academic Matters: The advisor and student discuss an academic topic of immediate relevance to their degree, including scheduling, academic administrative issues like degree requirements, major/minor selection, academic performance, or review of institutional benchmarks.
    “Student is really liking three of their four classes. MATH [course number] is giving them a big difficulty, and they know they cannot continue. They did do Summer Success292929a program for incoming students to take summer credits, so they will still be on track for a four year plan. Student likes summer, and planned to take math again in the summer. Student needs to talk to academic coach to release early alert grade.”
  • Non-academic Matters: The advisor and the student discuss a non-academic topic. This category includes career aspirations, planning for graduate school, financial matters such as aid or non-payment of tuition, extracurriculars, and personal life. While the discussion may touch on academics (for example, discussing academic context in the context of deciding a career or graduate school), it must go beyond the scope of immediate discussion of the student’s course performance or plans in their undergraduate degree.
    “Student did Summer Success, so they have made friends. But they are currently living with friends and family. [Redacted], because of such, do not have a permanent place to call home.”
  • Relationship Development: The advisor demonstrates development of a relationship or emotional connection to the student. This can come in the form of expressing concern, understanding, or support for students, noting the student’s emotional state during the interaction, or recording observations about the student’s personality.
    “New freshman — an introvert, so while they didn’t have too much to say how their classes or GSU was going, I could tell subtly that they did like it.”
B.2 Results
To supplement our analysis of results in §6.3, we report the frequency of codes by each advisor in Table 15.
Advisor 1 Advisor 2 Advisor 3
Code (851 comments) (426 comments) (375 comments)
Academic 654 343 290
Non=academic 153 63 135
Relationship 63 17 101
@@ B.2 Results
Table 15: Frequencies of the applied codes by each advisor during student-advising meetings, based on our exploratory qualitative analysis. Code counts per advisor may sum to a different number than the total number of comments per advisor as meeting comments can have zero or multiple codes applied.
Academic vs. holistic style
Here we expand on what we call the “academic style” of Advisor 1 and 2. There are two features of the academic style of advising. First, it tends to be very structured, especially for Advisor 1’s comments, which often begin with a routinized list of advising todos, followed by their proposed schedule:
“Student came to the office to discuss schedule for spring 2018 semester, printed and reviewed degree audit, student was shown major map, reviewed EAB for success markers, and notified student. Proposed schedule is: [courses]. Additional comments: [any additional comments].”
Out of 851 comments left by Advisor 1, 491 begin with “student came to the office to discuss” and follow a structure identical or quite similar to the one above. Interestingly, we notice that Advisor 1 is the only advisor who refers routinely to the use of digital interfaces (“reviewed EAB for success markers”) during advising.
Second, the focus is strictly academic. When Advisors 1 and 2 do personalize their advice to the student, their scope of consideration is also largely constrained to academic topics like placement or exams. For example, Advisor 2 writes:
“Student wants to change major from bio to exercise science… I advised that they could possibly do mini-mester for spring in order to compensate for credits lost towards degree.”
This difference in structure and focus on academic topics is reflected in the average Jaccard scores and overall distribution of codes.
B.3 Qualitative Analysis of Advisor Expertise
B.3.1 Codebook for Qualitative Analysis of Non-Algorithmic Information
@@
B.3.2 Comments Mentioning Reasons
@@
Instead of analyzing reasons for taking the intervention for mentions of non-algorithmic information, we analyzed comments for customizations of schedules. See Appendix B.3.3 for details about schedule customization.
B.3.3 Results for Intervention 17 (Recommend Schedule)
@@
B.3.4 Comparison to Insignificant Interventions
Intervention 5 Intervention 9 Intervention 10
Adjust Course Balance Address Financial Issue Address Non-Financial Requirement
1 Course Credits and Placement 5 - -
2 Course Problem 6 - -
3 Degree Progress 3 - -
4 Documents - 5 -
5 Fin. Registration Problem - 52 -
6 Load Management 1 - -
7 Major Choice 1 - -
8 Non-fin. Registration Problem - - 3
9 Personal Circumstances - 1 -
10 Scholarship Process - 6 -
11 Student-Initiated - 1 -
Table 17: Reasons cited by advisors for taking ExpertTest insignificant interventions.
Results for the insignificant interventions are reported in Table 17. Unlike the significant interventions, which incorporate diverse, non-algorithmic information, two of the three insignificant interventions, Intervention 9 (Address Financial Issues) and Intervention 10 (Address Non-financial Collegiate Requirements), are dominated by administrative holds on registration, most commonly due to nonpayment of a balance on the student’s account. These administrative holds are plausibly either predictable from administrative data (for instance, financial holds might be predictable from students’ expected financial contribution) or not outcome-relevant (when the holds are resolved).
@@ 11 Student-Initiated - 1 -
Appendix C Advisor Heterogeneity Analysis
C.1 Heterogeneity Analysis Methods
To examine heterogeneity in treatment effects by advisor assignment, we estimate linear and logistic regression models predicting graduation outcomes with interaction terms for the most common MAAPS advisors.
C.1.1 Data Cleaning
We perform the exclusions described in 3.1.3, resulting in a final test population of $n=834$ students ($427$ in treatment and $407$ in control).
We then define the following variables:
  • $T_{i}\in\{0,1\}$ indicates assignment to the treatment group
  • $Z_{i}\in\{0,1\}$, indicates that student $i$ graduated by 2020
  • $A_{ij}\in\{0,1\}$ for $j\in[1,3]$, where $A_{ij}=1$ indicating student $i$ was assigned to MAAPS advisor jj. [^30] For all students in the control arm, $A_{ij}=0$ for all $j$. When students in the MAAPS treatment arm met with multiple advisors, they are assigned the advisor they met with the most. There are five MAAPS advisors; advisors 4 and 5 are left out of the regression because they did not leave many comments on student meetings, and to avoid multicollinearity.
  • $\textsf{Gender}\in\{0,1\}$, with 0 indicating female and $1$ indicating male.
  • $\textsf{ACT}_{i}\in[0,36]$, the student’s ACT score. Following the methodology in Rossman et al. [77], we convert composite SAT scores to ACT scores using a concordance table provided by the Educational Testing Service and ACT Inc. [1].
  • $\textsf{EFC}_{i}\in[1,14]$ a scale indicating the level of the student’s expected financial contribution (from $\mathdollar 0-\mathdollar 1,050$ at level 0 to $\mathdollar 22,051+$ at level $14$).
  • $\textsf{Hours}_{i}\in\mathbb{Z}$ indicating the number of college credit hours earned by the student before enrollment at GSU, including AP or IB credit.
C.1.2 Regression
To examine heterogeneity in treatment effects by advisor assignment, we estimate a linear regression model predicting the binary graduation outcome:
$\mathbb{P}(Z_{i}=1)=$ $\beta_{0}+\beta_{1}T_{i}+\sum_{j}\theta_{j}A_{ij}+\gamma_{1}\textsf{Gender}_{i}+\gamma_{2}\textsf{ACT}_{i}+\gamma_{3}\textsf{EFC}_{i}+\gamma_{4}\textsf{Hours}_{i}$
We estimate an analogous logistic regression model to ensure consistency of advisor coefficient rankings.
C.2 Heterogeneity Analysis Results
While we do not find significant heterogeneous treatment effects across advisors, both linear (Table 18) and logistic regression (Table 19) produce consistent ranking of advisor-specific coefficients — with Advisor $3$ having the highest coefficient ($0.0685$ in the linear regression), followed by Advisor 2, and last Advisor 1.
OLS Regression Results: Advisor Heterogeneity in Graduation
Variable Coefficient Standard error $t$-statistic pp-value 95% CI
$\beta_{0}$ 0.2857 0.024 11.945 0.000 [0.239, 0.333]
$T_{i}$ 0.0573 0.082 0.700 0.484 [-0.103, 0.218]
$A_{i}^{1}$ -0.0833 0.088 -0.945 0.345 [-0.256, 0.090]
$A_{i}^{2}$ -0.0125 0.087 -0.144 0.886 [-0.183, 0.158]
$A_{i}^{3}$ 0.0685 0.087 0.785 0.433 [-0.103, 0.240]
Gender -0.1429 0.032 -4.524 0.000 [-0.205, -0.081]
ACT 0.0456 0.016 2.811 0.005 [0.014, 0.078]
EFC -0.0082 0.015 -0.550 0.583 [-0.038, 0.021]
Hours Other 0.0200 0.016 1.259 0.208 [-0.011, 0.051]
Observations: 834, Adjusted R-squared: 0.036, F-statistic: 4.889, F-statistic $p$-value: $6.48\times 10^{-6}$
Table 18: OLS regression estimating the effect of student assignment to a particular advisor on graduation outcomes. The model includes controls for gender, ACT scores, expected family contribution (EFC), and other academic hours applied toward the degree. Coefficient estimates, associated $t$-statistics, $p$-values, and 95% confidence intervals are reported.
Logistic Regression Results: Advisor Heterogeneity in Graduation
Variable Coefficient Standard error $z$-statistic pp-value 95% CI
$\beta_{0}$ -0.9454 0.127 -7.419 0.000 [-1.195, -0.696]
$T_{i}$ 0.3265 0.438 0.745 0.456 [-0.533, 1.186]
$A_{i}^{1}$ -0.4799 0.479 -1.002 0.316 [-1.419, 0.459]
$A_{i}^{2}$ -0.0877 0.465 -0.188 0.851 [-0.999, 0.824]
$A_{i}^{3}$ 0.3160 0.460 0.687 0.492 [-0.586, 1.218]
Gender -0.8024 0.182 -4.404 0.000 [-1.160, -0.445]
ACT 0.2466 0.088 2.879 0.005 [0.073, 0.420]
EFC -0.0407 0.082 -0.496 0.620 [-0.202, 0.120]
Hours Other 0.0902 0.082 1.098 0.272 [-0.071, 0.251]
Observations: 834, Log likelihood: -458.97, Log likelihood null: -477.05, LLR $p$-value: $7.044\times 10^{-6}$
Table 19: Logistic regression estimating the effect of student assignment to a particular advisor on graduation outcomes. The model includes controls for gender, ACT scores, expected family contribution (EFC), and other academic hours applied toward the degree. Coefficient estimates, associated $z$-statistics, $p$-values, and 95% confidence intervals are reported.